### v1

@@ Abstract
keywords:
Video Text Spotting, Video Text Dataset, Tracking-by-Detection, Parameter- and Data-Efficient
1 Introduction
@@
Figure 1: (a) ‘Gap between Spot. & Det.’: the gap between spotting and detection F1-score. As the spotting task involves recognizing the results of the detection process, the detection score is indeed the upper bound of spotting performance. The larger the gap, the poorer the recognition ability. Compared to the ITS model (Deepsolo [37]), the VTS model (TransDETR [29]) presents unsatisfactory image-level text spotting F1-scores, which lag far behind its detection performance, especially on ArTVideo with curved text. It indicates recognition capability is a main bottleneck in the VTS model. (b) GoMatching++ outperforms TransDETR by over 9.68 MOTA on ICDAR15-video while saving 301 training GPU hours and 19.4GB memory.
@@
With the burgeoning growth of visual information technology, text spotting has emerged as an effective tool for various applications, such as video understanding [27], video retrieval [8] and autonomous driving [43, 38]. Recently, numerous image text spotting (ITS) methods [20, 9, 25, 42, 37, 13] that simultaneously tackle text detection and recognition, have attained extraordinary accomplishment. In the video realm, video text spotting (VTS) introduces an additional tracking task, intensifying the challenge to a new level. Although VTS methods [30, 4, 5, 33, 31, 29] have made significant progress, a substantial discrepancy persists when compared to ITS. We observe that the text recognition proficiency of VTS models is far inferior to ITS models, even when extensively trained end-to-end on video data.
@@
To investigate this disparity, we compare the state-of-the-art (SOTA) VTS model TransDETR [29] and ITS model Deepsolo [37] for image-level text spotting performance on ICDAR15-video [14] and our ArTVideo (i.e., Arbitrary-shaped Text in Video) (Sec.4), which comprises over 30% curved text. As illustrated in Fig. 1(a), even when evaluating the image-level spotting performance on the VTS model’s training set, the F1-score of TransDETR is merely comparable to the zero-shot performance of Deepsolo. The performance of the VTS model on ArTVideo is much worse. Moreover, the significant disparity between the spotting and detection-only performance of the VTS model highlights that its recognition capability constitutes the main bottleneck. We attribute this discrepancy to two key aspects: 1) the model architecture and 2) the training data. First, in terms of model architecture, some studies [44, 35] have indicated that there exists optimization conflict in detection and association during the end-to-end training of MOTR [39]. We hold that TransDETR [29], which integrates text recognition into a MOTR-based design, is likely to incur more serious conflicts, alongside requiring extensive training resources and data in an end-to-end framework. On the other hand, ITS studies [37, 13] have presented the advantages of employing advanced query formulation for text spotting in DETR frameworks [6, 41]. In contrast, existing Transformer-based VTS models still rely on Region of Interest (RoI) components or simply cropping detected text regions for recognition. Second, with respect to the training data, most text instances in current video text dataset [14, 33, 34] are either horizontal or oriented, and the bounding box annotations are limited to quadrilaterals. This constrains both the data diversity and the recognition performance, particularly for curved text. Hence, leveraging model and data knowledge from ITS presents considerable value for VTS. To achieve this, an intuitive idea is to exploit an off-the-shelf SOTA image text spotter. However, the utilization in a naive manner gives rise to two fundamental problems: 1) video frames often feature a higher frequency of small and blurry text instances caused by camera motion and varying distances. This domain gap leads to low detection confidence and consequently a relatively low Recall for the image text spotter, and 2) native image text spotters lack the ability to track the text across frames.
@@
To this end, we propose GoMatching++, a simple and effective baseline that adheres to the tracking-by-detection paradigm. Specifically, building on a frozen ITS model (e.g.DeepSolo [37]), we introduced the rescoring mechanism along with a tailored tracker, LST-Matcher. The rescoring mechanism involves an additional lightweight, trainable text classification head called rescoring head via efficient tuning on video data and recalibrating confidence scores for detected instances. Each instance’s final score is computed through a fusion operation between the original score from the image text spotter and the recalibrated score produced by the rescoring head. This mechanism preserve the ITS model’s inherent knowledge while alleviating performance degradation arising from the image-video domain gap. Subsequently, the identified text instances are sent to LST-Matcher for association. LST-Matcher, consisting of a shared feedforward network (FFN) and dual parallel Transformers with single-layer encoder-decoder architectures, can effectively harness both long- and short-term information, endowing it a high-capable tracker. Since both the rescoring head and LST-Matcher are lightweight and gradient-decoupled, GoMatching++ significantly reduces training resources and data requirements while avoiding optimization conflicts.
@@
This work is an extensive version of our conference paper [12], offering four additional contributions: (1) We explore more concise architectural designs for GoMatching and propose GoMatching++, a novel parameter- and data-efficient model. (2) We expand ArTVideo from a simple test set to a comprehensive benchmark, including both training and test sets, to mitigate the scarcity of curved text datasets in the VTS domain. (3) We conduct extensive experiments to ablate and analyze the efficiency and effectiveness of GoMatching++. (4) We provide an in-depth statistical and experimental analysis of ArTVideo, providing new insights into VTS.
@@
The rest of this paper is organized as follows. In Sec. 2, we briefly review the related work. In Sec. 3, we introduce our proposed method in detail. Sec. 4 provides a comprehensive overview of the benchmark. Sec. 5 and Sec. 6 report and discuss our experimental results on existing open-source datasets and our ArtVideo, respectively. Lastly, we conclude our study in Sec. 7.
2 Related Work
2.1 Image Text Spotting
@@
Early approaches [22, 32, 23] crafted RoI-based modules to bridge text detection and recognition. However, these methods ignored one vital issue, i.e., the synergy problem between the two tasks. To overcome this dilemma, recent Transformer-based methods [15, 25, 36, 13] get rid of the fetters of RoI modules, and chase a better representation for the two tasks. For example, DETR-based TESTR [42] uses two decoders for each task in parallel. In contrast, DeepSolo [37] proposes a unified and explicit query form for the two tasks, without harnessing dual decoders. However, the above methods cannot perform tracking in the video.
2.2 Video Text Spotting
@@
Compared to ITS, existing SOTA VTS methods still rely on RoI or simply cropping detected text regions for recognition. TransVTSpotter [33] directly crops the detection regions from a transformer-based detector to perform recognition. CoText [31] adopts a text spotter with Masked-RoI and applies cosine similarity matching for tracking via a tracking head. TransDETR [29] performs detection and tracking under the MOTR paradigm, using Rotated-RoI for the subsequent recognizer. They pursue training all modules in an end-to-end manner. In comparison, we explore turning a RoI-free ITS model into a VTS one. We reveal the probability of freezing ITS part and focusing on tracking, thereby establishing a novel parameter- and data-efficient baseline while achieving SOTA performance.
2.3 Video Text Datasets
Current progress in video text spotting research can also be attributed to the publicly accessible datasets, which serve as critical benchmarks for evaluating and advancing methods in the field. ICDAR2015-video [14] contains 49 video clips, where the locations of words in the video are labeled using oriented bounding boxes. BOVText [33] introduces a large-scale, bilingual (English and Chinese), open-world video text dataset, providing 2000+ videos with various scenarios. Recently, DSText [34] establishes a dense and small text video dataset, including 100 videos. However, these datasets predominantly comprise straight text with almost no curved text, and their annotation boxes are restricted to quadrilateral shapes. To advance research in video curved text spotting, we introduce a novel benchmark comprising 60 videos, with over 30% featuring curved text instances, designed for both training and test.
@@ 2.3 Video Text Datasets
3 Methodology
3.1 Overview
@@ 3.2 Rescoring Mechanism
3.3 Long-Short-Term Matching Module
@@
where $S_{ij}$ represents the similarity score derived from cosine similarity. Then, a cross-entropy loss function is used to optimize the predicted distributions following GTR [45].
During inference, we engage a memory bank to store the instance trajectories from $H$ history frames for long term association. All detected text instances in each frame are further processed by non-maximum-suppression (NMS) before being fed into LST-Matcher for association. Unlike in the training phase, where ST-Matcher and LT-Matcheder are independent of each other, LST-Matcher comprises a two-stage association procedure as described in Fig. 3. Concretely, ST-Matcher first matches the embedding $E^{t}$ in the current frame $t$ with the trajectories $\tau^{t-1}$ in the previous frame $t-1$. Then, LT-Matcher employs other trajectories $\tau^{H}_{oth}$ in the memory bank to associate the unmatched ones $E^{t}_{s\_u}$ with low association score in ST-Matcher caused by missed detections or strong appearance changes. If the association score with any trajectory calculated in ST-Matcher or LT-Matcher is higher than a threshold $\theta$, the instance is linked to the trajectory with the highest score. Otherwise, this instance is used to initiate a new trajectory. Finally, we combine the trajectories $\tau_{s}$ and $\tau_{l}$ predicted by ST-Matcher and LT-Matcher to obtain new trajectories $\tau^{N}$ for tracking in the next frame.
3.4 Parameter- and Data-Efficient Design
@@
(a) Transformer-based: The ST-Matcher and LT-Matcher both incorporate a single-layer Transformer encoder and a single-layer Transformer decoder, sharing the FFN with two linear layers (denoted by ‘*’). This is the exact design of the GoMatching model as presented in the conference version [12]. The instance queries $Q^{t}$ from current frame $t$ are first projected into a new feature space by the FFN and obtain the embeddings $E^{t}$. In the encoder, the historical embeddings $E^{H}$ are enhanced by self-attention. The decoder takes embeddings in the current frame as query and enhanced historical embeddings as key for cross-attention, and computes the association score matrix $G$.
(b) Similarity-only: The instance queries $Q^{t}$ and $Q^{H}$ are used to directly compute the association score matrix $G$. However, the absence of the FFN and attention mechanism to map queries into a more suitable feature space for similarity calculation leads to a significant degradation in performance (Sec. 5.3).
(c) FFN-based: The instance queries $Q^{t}$ are first projected by the FFN to derive the embeddings $E^{t}$, which are then used to compute the association score matrix $G$ with the historical embeddings $E^{H}$. Similarly, The absence of an attention mechanism to focus on key instance features leads to a performance decline (Sec. 5.3).
(d) Cross-attn-based: The ST-Matcher and LT-Matcher are designed with a more elegant structure to obtain $G$, incorporating shared FFN and cross-attention. Notably, the number of parameters of this structure is approximately $1/3$ that of the Transformer-based architecture (a), while maintaining comparable or even superior performance (Sec. 5.3).
By freezing the image text spotter and adopting the lightweight design for the proposed modules, GoMatching++ significantly reduces the number of trainable parameters, establishing a more parameter-efficient baseline. Additionally, the decoupled framework eliminates the need for joint optimization of detection, recognition, and tracking, enabling the matcher to fully leverage the learned features from the image text spotter. This design reduces computational overhead, stabilizes training, and enhances the use of limited data while achieving high tracking performance. Further experimental details are provided in Sec. 5.3 and Sec. 5.4.
3.5 Optimization
Rescoring Loss. To train the rescoring head, we following DETR [6] and use Hungarian algorithm [17] to find a bipartite matching $\hat{\sigma}$ between the predictions $\hat{Y}$ and the ground truths $Y$ with minimum matching cost $\mathcal{C}$:
@@ 3.5 Optimization
where $\lambda_{c}$ and $\lambda_{b}$ serve as the hyper-parameters to balance different tasks. $\hat{p}_{\sigma(i)}(c_{i})$ and $\hat{b}_{\sigma(i)}$ are the probability for ground truth class $c_{i}$ and the predicition of bounding box respectively, and $b_{i}$ represents the ground truth bounding box. $\mathcal{L}_{cls}$ is the focal loss [18]. Specifically, the focal loss for training the rescoring head can be formulated as:
@@
Long-Short Association Loss. In ST-Matcher, only trajectories within adjacent frames of the input batch are considered, whereas LT-Matcher extends its scope to trajectories across all frames within the batch, using long-term temporal information. For each trajectory, we optimize the log-likelihood of its assignments $\hat{\alpha}_{k}$ by cross-entropy, following GTR [45]:
@@
Figure 4: Visual examples from ArTVideo. The straight and curved text are labeled with quadrilaterals and polygons, respectively. The same background color in different frames denotes the same instance.
@@
ICDAR15-video[14] 49 27k 144k straight quadrilateral $\times$
BOVText[33] 2,021 1.7m 8.8m straight quadrilateral $\times$
DSText[34] 100 56k 671k straight quadrilateral $\times$
@@
Table 1: Comparison with existing video text datasets. ‘Text’ and ‘Box’ denote the types of text instances and annotation formats, respectively, while ‘Mask’ indicates whether each text instance includes mask annotations.

### v2

@@ Abstract
I Introduction
((a))
((b))
@@ I Introduction
Figure 1: (a) ‘Gap between Spot. & Det.’: the gap between spotting and detection F1-score. As the spotting task involves recognizing the results of the detection process, the detection score is indeed the upper bound of spotting performance. The larger the gap, the poorer the recognition ability. Compared to the ITS model (Deepsolo [ye2023deepsolo]), the VTS model (TransDETR [wu2024end]) presents unsatisfactory image-level text spotting F1-scores, which lag far behind its detection performance, especially on ArTVideo with curved text. It indicates recognition capability is a main bottleneck in the VTS model. (b) GoMatching++ outperforms TransDETR by over 9.68 MOTA on ICDAR15-video while saving 301 training GPU hours and 19.4GB memory.
@@
With the burgeoning growth of visual information technology, text spotting has emerged as an effective tool for various applications, such as video understanding [srivastava2015unsupervised, zhao2022towards, tom2023reading], video retrieval [dong2021dual] and autonomous driving [zhang2021character, zablocki2022explainability]. Recently, numerous image text spotting (ITS) methods [long2021scene, feng2021residual, liu2023spts, zhang2022text, ye2023deepsolo, huang2023estextspotter] that simultaneously tackle text detection and recognition, have attained extraordinary accomplishment. In the video realm, video text spotting (VTS) introduces an additional tracking task, intensifying the challenge to a new level. Although VTS methods [wang2017end, cheng2019you, cheng2020free, wu2021bilingual, wu2024bilingual, wu2024end] have made significant progress, a substantial discrepancy persists when compared to ITS. We observe that the text recognition proficiency of VTS models is far inferior to ITS models, even when extensively trained end-to-end on video data.
@@
To investigate this disparity, we compare the state-of-the-art (SOTA) VTS model TransDETR [wu2024end] and ITS model Deepsolo [ye2023deepsolo] for image-level text spotting performance on ICDAR15-video [Karatzas2015ICDAR15] and our ArTVideo (i.e., Arbitrary-shaped Text in Video) (Sec.IV), which comprises over 30% curved text. As illustrated in Fig. 1(a), even when evaluating the image-level spotting performance on the VTS model’s training set, the F1-score of TransDETR is merely comparable to the zero-shot performance of Deepsolo. The performance of the VTS model on ArTVideo is much worse. Moreover, the significant disparity between the spotting and detection-only performance of the VTS model highlights that its recognition capability constitutes the main bottleneck. We attribute this discrepancy to two key aspects: 1) the model architecture and 2) the training data. First, in terms of model architecture, some studies [zhang2023motrv2, yu2023motrv3] have indicated that there exists optimization conflict in detection and association during the end-to-end training of MOTR [zeng2022motr]. We hold that TransDETR [wu2024end], which integrates text recognition into a MOTR-based design, is likely to incur more serious conflicts, alongside requiring extensive training resources and data in an end-to-end framework. On the other hand, ITS studies [ye2023deepsolo, huang2023estextspotter] have presented the advantages of employing advanced query formulation for text spotting in DETR frameworks [carion2020end, zhu2020deformable]. In contrast, existing Transformer-based VTS models still rely on Region of Interest (RoI) components or simply cropping detected text regions for recognition. Second, with respect to the training data, most text instances in current video text dataset [Karatzas2015ICDAR15, wu2021bilingual, Wu2023DSText] are either horizontal or oriented, and the bounding box annotations are limited to quadrilaterals. This constrains both the data diversity and the recognition performance, particularly for curved text. Hence, leveraging model and data knowledge from ITS presents considerable value for VTS. To achieve this, an intuitive idea is to exploit an off-the-shelf SOTA image text spotter. However, the utilization in a naive manner gives rise to two fundamental problems: 1) video frames often feature a higher frequency of small and blurry text instances caused by camera motion and varying distances. This domain gap leads to low detection confidence and consequently a relatively low recall for the image text spotter, and 2) native image text spotters lack the ability to track the text across frames.
@@
To this end, we propose GoMatching++, a simple and effective baseline that adheres to the tracking-by-detection paradigm. Specifically, building on a frozen ITS model (e.g.DeepSolo [ye2023deepsolo]), we introduced the rescoring mechanism along with a tailored tracker, LST-Matcher. The rescoring mechanism involves an additional lightweight, trainable text classification head called rescoring head via efficient tuning on video data and recalibrating confidence scores for detected instances. Each instance’s final score is computed through a fusion operation between the original score from the image text spotter and the recalibrated score produced by the rescoring head. This mechanism preserve the ITS model’s inherent knowledge while alleviating performance degradation arising from the image-video domain gap. Subsequently, the identified text instances are sent to LST-Matcher for association. LST-Matcher, consisting of a shared feedforward network (FFN) and dual parallel Transformers with single-layer encoder-decoder architectures, can effectively harness both long- and short-term information, endowing it a high-capable tracker. Since both the rescoring head and LST-Matcher are lightweight and gradient-decoupled, GoMatching++ significantly reduces training resources and data requirements while avoiding optimization conflicts.
@@
This work is an extensive version of our conference paper [he2024gomatching], offering four additional contributions: (1) We explore more concise architectural designs for GoMatching and propose GoMatching++, a novel parameter- and data-efficient model. (2) We expand ArTVideo from a simple test set to a comprehensive benchmark, including both training and test sets, to mitigate the scarcity of curved text datasets in the VTS domain. (3) We conduct extensive experiments to ablate and analyze the efficiency and effectiveness of GoMatching++. (4) We provide an in-depth statistical and experimental analysis of ArTVideo, providing new insights into VTS.
@@
The rest of this paper is organized as follows. In Sec. II, we briefly review the related work. In Sec. III, we introduce our proposed method in detail. Sec. IV provides a comprehensive overview of the benchmark. Sec. V and Sec. VI report and discuss our experimental results on existing open-source datasets and our ArtVideo, respectively. Lastly, we conclude our study in Sec. VII.
II Related Work
II-A Image Text Spotting
@@
Early approaches [liao2020mask, wang2021pan++, liu2021abcnet] crafted RoI-based modules to bridge text detection and recognition. However, these methods ignored one vital issue, i.e., the synergy problem between the two tasks. To overcome this dilemma, recent Transformer-based methods [cao2021all, zhang2022text, liu2023spts, ye2023deepsolo, tong2024granularity] get rid of the fetters of RoI modules, and chase a better representation for the two tasks. For example, DETR-based TESTR [zhang2022text] uses two decoders for each task in parallel. In contrast, DeepSolo [ye2023deepsolo] proposes a unified and explicit query form for the two tasks, without harnessing dual decoders. However, the above methods cannot perform tracking in the video.
II-B Video Text Spotting
@@
Compared to ITS, existing SOTA VTS methods still rely on RoI or simply cropping detected text regions for recognition. TransVTSpotter [wu2021bilingual] directly crops the detection regions from a transformer-based detector to perform recognition. CoText [wu2024bilingual] adopts a text spotter with Masked-RoI and applies cosine similarity matching for tracking via a tracking head. TransDETR [wu2024end] performs detection and tracking under the MOTR paradigm, using Rotated-RoI for the subsequent recognizer. They pursue training all modules in an end-to-end manner. In comparison, we explore turning a RoI-free ITS model into a VTS one. We reveal the probability of freezing ITS part and focusing on tracking, thereby establishing a novel parameter- and data-efficient baseline while achieving SOTA performance.
@@
II-C Video Text Datasets
Current progress in video text spotting research can also be attributed to the publicly accessible datasets, which serve as critical benchmarks for evaluating and advancing methods in the field. ICDAR2015-video [Karatzas2015ICDAR15] contains 49 video clips, where the locations of words in the video are labeled using oriented bounding boxes. BOVText [wu2021bilingual] introduces a large-scale, bilingual (English and Chinese), open-world video text dataset, providing 2000+ videos with various scenarios. Recently, DSText [Wu2023DSText] establishes a dense and small text video dataset, including 100 videos. However, these datasets predominantly comprise straight text with almost no curved text, and their annotation boxes are restricted to quadrilateral shapes. To advance research in video curved text spotting, we introduce a novel benchmark comprising 60 videos, with over 30% featuring curved text instances, designed for both training and test.
III Methodology
III-A Overview
@@ III-B Rescoring Mechanism
III-C Long-Short-Term Matching Module
@@
where $S_{ij}$ represents the similarity score derived from cosine similarity. Then, a cross-entropy loss function is used to optimize the predicted distributions following GTR [zhou2022global].
During inference, we engage a memory bank to store the instance trajectories from $H$ history frames for long term association. All detected text instances in each frame are further processed by non-maximum-suppression (NMS) before being fed into LST-Matcher for association. Unlike in the training phase, where ST-Matcher and LT-Matcheder are independent of each other, LST-Matcher comprises a two-stage association procedure as described in Fig. 3. Concretely, ST-Matcher first matches the embedding $E^{t}$ in the current frame $t$ with the trajectories $\tau^{t-1}$ in the previous frame t1t-1. Then, LT-Matcher employs other trajectories $\tau^{H}_{oth}$ in the memory bank to associate the unmatched ones $E^{t}_{s\_u}$ with low association score in ST-Matcher caused by missed detections or strong appearance changes. If the association score with any trajectory calculated in ST-Matcher or LT-Matcher is higher than a threshold $\theta$, the instance is linked to the trajectory with the highest score. Otherwise, this instance is used to initiate a new trajectory. Finally, we combine the trajectories $\tau_{s}$ and $\tau_{l}$ predicted by ST-Matcher and LT-Matcher to obtain new trajectories $\tau^{N}$ for tracking in the next frame.
III-D Parameter- and Data-Efficient Design
@@
(a) Transformer-based: The ST-Matcher and LT-Matcher both incorporate a single-layer Transformer encoder and a single-layer Transformer decoder, sharing the FFN with two linear layers (denoted by ‘*’). This is the exact design of the GoMatching model as presented in the conference version [he2024gomatching]. The instance queries $Q^{t}$ from current frame $t$ are first projected into a new feature space by the FFN and obtain the embeddings $E^{t}$. In the encoder, the historical embeddings $E^{H}$ are enhanced by self-attention. The decoder takes embeddings in the current frame as query and enhanced historical embeddings as key for cross-attention, and computes the association score matrix $G$.
(b) Similarity-only: The instance queries $Q^{t}$ and $Q^{H}$ are used to directly compute the association score matrix $G$. However, the absence of the FFN and attention mechanism to map queries into a more suitable feature space for similarity calculation leads to a significant degradation in performance (Sec. V-C).
(c) FFN-based: The instance queries $Q^{t}$ are first projected by the FFN to derive the embeddings $E^{t}$, which are then used to compute the association score matrix $G$ with the historical embeddings $E^{H}$. Similarly, The absence of an attention mechanism to focus on key instance features leads to a performance decline (Sec. V-C).
(d) Cross-attn-based: The ST-Matcher and LT-Matcher are designed with a more elegant structure to obtain $G$, incorporating shared FFN and cross-attention. Notably, the number of parameters of this structure is approximately $1/3$ that of the Transformer-based architecture (a), while maintaining comparable or even superior performance (Sec. V-C).
By freezing the image text spotter and adopting the lightweight design for the proposed modules, GoMatching++ significantly reduces the number of trainable parameters, establishing a more parameter-efficient baseline. Additionally, the decoupled framework eliminates the need for joint optimization of detection, recognition, and tracking, enabling the matcher to fully leverage the learned features from the image text spotter. This design reduces computational overhead, stabilizes training, and enhances the use of limited data while achieving high tracking performance. Further experimental details are provided in Sec. V-C and Sec. V-D.
III-E Optimization
Rescoring Loss. To train the rescoring head, we following DETR [carion2020end] and use Hungarian algorithm [kuhn1955hungarian] to find a bipartite matching $\hat{\sigma}$ between the predictions $\hat{Y}$ and the ground truths $Y$ with minimum matching cost $\mathcal{C}$:
@@ III-E Optimization
where $\lambda_{c}$ and $\lambda_{b}$ serve as the hyper-parameters to balance different tasks. $\hat{p}_{\sigma(i)}(c_{i})$ and $\hat{b}_{\sigma(i)}$ are the probability for ground truth class $c_{i}$ and the predicition of bounding box respectively, and $b_{i}$ represents the ground truth bounding box. $\mathcal{L}_{cls}$ is the focal loss [lin2017focal]. Specifically, the focal loss for training the rescoring head can be formulated as:
@@
Long-Short Association Loss. In ST-Matcher, only trajectories within adjacent frames of the input batch are considered, whereas LT-Matcher extends its scope to trajectories across all frames within the batch, using long-term temporal information. For each trajectory, we optimize the log-likelihood of its assignments $\hat{\alpha}_{k}$ by cross-entropy, following GTR [zhou2022global]:
@@
Figure 4: Visual examples from ArTVideo training set. The straight and curved text are labeled with quadrilaterals and polygons, respectively. The same background color in different frames denotes the same instance.
@@
ICDAR15-video[Karatzas2015ICDAR15] 49 27k 144k straight quadrilateral $\times$
BOVText[wu2021bilingual] 2,021 1.7m 8.8m straight quadrilateral $\times$
DSText[Wu2023DSText] 100 56k 671k straight quadrilateral $\times$
@@
TABLE I: Comparison with existing video text datasets. ‘Text’ and ‘Box’ denote the types of text instances and annotation formats, respectively, while ‘Mask’ indicates whether each text instance includes mask annotations.
IV The ArTVideo Benchmark
IV-A Data Source and Annotation